OpenStudy (anonymous):

please help! .M is the midpoint of AB. The coordinates of A are (2,3) and the coordinates of M are (4,5,6) find the coordinates of B/

OpenStudy (anonymous):

can u solve for y,z now

OpenStudy (anonymous):

multiple choices are: a. (3.25,4.5) b.(7,9) c. (2.5),3) d. (11, 15)

OpenStudy (anonymous):

what r 2,3 x,y coordinates or y,z coordinates

OpenStudy (anonymous):

its correct x-2=8 x=10 y-3=10 y=13

OpenStudy (anonymous):

so which one of thos multiple choice would it be?

OpenStudy (anonymous):

check that these choices are given

OpenStudy (anonymous):

it would be D right?

OpenStudy (anonymous):

may be choices are given wrong its not d

OpenStudy (anonymous):

no, this is for school. they gave me these multiple choiceds

OpenStudy (anonymous):

I need to go to bed, can u just tell me what it is?

OpenStudy (anonymous):

may be you have wrriten wrong or your teacher gave wrong opt to confuse you show my solution to your teacher and message me when u get answer

OpenStudy (anonymous):

nvm...

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

by distance fomula d=x2-x1 f0r mid point d=(x2-x1)/2